# literature.bib

### v4

@@ 4.1. Weak convergence approach
if and only if $Y^{\varepsilon}$ is a strong solution in the sense of [AV22variational, Def. 3.2] to (4.1) with $x$ replaced by $\xi$. This is a mere consequence of (A.7) and the fact that $B(\cdot,Y(\cdot))\in L^{2}(([0,T]\times\Omega,\mathcal{P},\lambda\times\mathbb{P});\mathcal{L}_{2}(U,H))\subset\mathcal{N}(0,T)$ for any $Y\in\mathrm{MR}(0,T)$. By [AV22variational, Th. 3.5], (4.1) has a unique strong solution $Y^{\varepsilon}$, also when $x$ is replaced by random initial data $\xi$. Thus we have pathwise uniqueness in the sense of [rock08, Def. 1.7] and we have existence of a.s. $\mathrm{MR}(0,T)$-valued weak solutions. Now fix $x\in H$ and $\varepsilon>0$ and let $Y^{\varepsilon}$ be the unique strong solution to (4.1). By [rock08, Th. 2.1, Def. 1.9(2), Def. 1.8] there exists a measurable map $\mathcal{G}^{\varepsilon}\colon C([0,T];U_{1})\to\mathrm{MR}(0,T)$ such that a.s. $Y^{\varepsilon}=\mathcal{G}^{\varepsilon}(\tilde{W}_{1}(\cdot))$.
@@
We have $\mathbb{E}[\exp(\frac{1}{2}\|-\frac{1}{\sqrt{\varepsilon}}\Psi^{\varepsilon}\|_{L^{2}(0,T;U)}^{2})]\leq\exp(\frac{K^{2}}{2\varepsilon})<\infty$, so by Novikov’s condition [karatzas98, Prop. 5.12],
@@
Now Girsanov’s theorem [liurockner15, Proposition I.0.6], [ferrario13, Th. 2.3] yields that $\hat{W}$ is an ${\mathbb{R}}^{\infty}$-Brownian motion on $(\Omega,\mathcal{F},\hat{\mathbb{P}},(\mathcal{F}_{t})_{t\geq 0})$, where
^$\hat{\mathbb{P}}\coloneqq\exp\left(-\frac{1}{\sqrt{\varepsilon}}\int_{0}^{T}\langle\Psi^{\varepsilon}(s),\,\mathrm{d}\tilde{W}(s)\rangle_{U}-\frac{1}{2\varepsilon}\|\Psi^{\varepsilon}\|_{L^{2}(0,T;U)}^{2}\right)\,\mathrm{d}\mathbb{P}.$
@@
$\mathbb{P}$-a.s. in $U_{1}$, where we used that $\Psi^{\varepsilon}\in\mathcal{A}_{K}$ to apply Fubini’s theorem in the second line. Thus, recalling the definition of ${X}^{\varepsilon}$ and noting that $\hat{\mathbb{P}}\ll\mathbb{P}\ll\hat{\mathbb{P}}$, we have $\hat{\mathbb{P}}$-a.s. $X^{\varepsilon}=\mathcal{G}^{\varepsilon}(\hat{W}_{1}(\cdot))$. By the Yamada-Watanabe theorem [rock08, Th. 2.1, Def. 1.9], for ${X}^{\varepsilon}=\mathcal{G}^{\varepsilon}(\hat{W}_{1}(\cdot))$ we have that $({X}^{\varepsilon},\hat{W})$ is a weak solution to (4.1). That is, ${X}^{\varepsilon}$ satisfies $\hat{\mathbb{P}}$-a.s. in $V^{*}$:
@@
where the last equality follows from the definition of $\hat{\beta}_{k}$ and (A.6). Let $\hat{\mathcal{N}}(0,T)$ denote the stochastically integrable processes with respect to $\hat{\mathcal{W}}$ and $\hat{W}$ on $(\Omega,\mathcal{F},\hat{\mathbb{P}},(\mathcal{F}_{t})_{t\geq 0})$, i.e. (A.3) with $\mathbb{P}$ replaced by $\hat{\mathbb{P}}$. Note that $\hat{\mathcal{N}}(0,T)=\mathcal{N}(0,T)$, since $\mathbb{P}\ll\hat{\mathbb{P}}\ll\mathbb{P}$. Thus, Proposition A.4 gives $\int_{0}^{t}\Phi(s)\,\mathrm{d}\hat{\mathcal{W}}(s)=\int_{0}^{t}\Phi(s)\,\mathrm{d}\hat{W}(s)$$\hat{\mathbb{P}}$-a.s. for all $\Phi\in\mathcal{N}(0,T)$ and $t\in[0,T]$. Therefore, combined with (4.6), ${X}^{\varepsilon}$ satisfies $\hat{\mathbb{P}}$-a.s. (hence $\mathbb{P}$-a.s.) in $V^{*}$:
@@
In the last line we used that $\int_{0}^{t}\Phi(s)\,\mathrm{d}\hat{\mathcal{W}}(s)=\int_{0}^{t}\Phi(s)\,\mathrm{d}{W}(s)+\frac{1}{\sqrt{\varepsilon}}\int_{0}^{t}\Phi(s)\Psi^{\varepsilon}(s)\,\mathrm{d}s$ for $\Phi\in\mathcal{N}(0,T)$ and $t\in[0,T]$. For $\Phi=\mathbbm{1}_{A\times(t_{1},t_{2}]}\otimes(u\otimes x)$ with $0\leq t_{1}<t_{2}\leq T$, $A\in\mathcal{F}_{t_{1}}$, $u\in U$, $x\in H$, the identity follows from (4.7) and the definition of the stochastic integral for elementary processes [NVW15, p. 305]. By linearity and continuity of the integrals and by a density argument and localization, the identity extends for $\Phi\in\mathcal{N}(0,T)$. This finishes the proof of the last claim of the lemma. ∎
@@ 4.2. Weakly continuous dependence in the skeleton equation
In this subsection we prove that condition (ii) of Theorem 4.5 is satisfied. This will be achieved in the upcoming Proposition 4.9. Its proof was inspired by [hongliliu21, Th. 3.2]. Using an additional approximation by Bochner-simple functions, we can omit the time(-Hölder) regularity assumptions on $B$ of [hongliliu21, (H5)].
@@
To estimate further, we use an argument inspired by [hongliliu21, Lem. 3.3]. Note that $w_{n}(0)=0$ and
@@
For any $t\in[\delta,T]$, we can apply the chain rule (A.2) to $v_{n}^{t,\delta}(\cdot)\coloneqq w_{n}(\cdot)-w_{n}(t_{\delta})=\int_{t_{\delta}}^{\cdot}\alpha_{n}(s)\,\mathrm{d}s$ on $[t_{\delta},T]$ and obtain for all $\tilde{t}\in[t_{\delta},T]$:
@@
$I_{1}^{n}(t)$ $\leq\int_{0}^{t}-\langle\bar{A_{0}}w_{n}(s),w_{n}(s)\rangle+|\!|\!|\bar{B_{0}}w_{n}(s)|\!|\!|_{H}\|\psi_{n}(s)\|_{U}\|w_{n}(s)\|_{H}\,\mathrm{d}s$
$\leq\int_{0}^{t}-\langle\bar{A_{0}}w_{n}(s),w_{n}(s)\rangle+\frac{1}{2}|\!|\!|\bar{B_{0}}w_{n}(s)|\!|\!|_{H}^{2}+\frac{1}{2}\|\psi_{n}(s)\|_{U}^{2}\|w_{n}(s)\|_{H}^{2}\,\mathrm{d}s$
@@
$I_{2}^{n}(t)$ $\leq\int_{0}^{t}\|f_{n}(s)\|_{V^{*}}\|w_{n}(s)\|_{V}+|\!|\!|g_{n}(s)|\!|\!|_{H}\|\psi_{n}(s)\|_{U}\|w_{n}(s)\|_{H}\,\mathrm{d}s$
(4.21) $\leq\int_{0}^{t}\frac{1}{\theta_{N,T}}\|f_{n}(s)\|_{V^{*}}^{2}+\frac{\theta_{N,T}}{4}\|w_{n}(s)\|_{V}^{2}+\frac{1}{2}|\!|\!|g_{n}(s)|\!|\!|_{H}^{2}+\frac{1}{2}\|\psi_{n}(s)\|_{U}^{2}\|w_{n}(s)\|_{H}^{2}\,\mathrm{d}s.$
@@
By (4.24), it remains to show that $\lim_{n\to\infty}\sup_{t\in[0,T]}|I_{3}^{n}(t)|=0$. We use Lemma 4.8. Note that $\sup_{n\in{\mathbb{N}}}\|w_{n}\|_{\mathrm{MR}(0,T)}\leq N$ by (4.19), so we only have to verify boundedness of $(\alpha_{n})\subset L^{1}(0,T;V^{*})$, where $\alpha_{n}\coloneqq-\bar{A_{0}}w_{n}+f_{n}+(\bar{B_{0}}w_{n}+g_{n})\psi_{n}+{b}(\psi_{n}-\psi)\in L^{2}(0,T;V^{*})+L^{1}(0,T;H)\subset L^{1}(0,T;V^{*})$. The last inclusion is continuous, so it suffices to prove boundedness of $(-\bar{A_{0}}w_{n}+f_{n})\subset L^{2}(0,T;V^{*})$ and $(\beta_{n})\coloneqq((\bar{B_{0}}w_{n}+g_{n})\psi_{n}+{b}(\psi_{n}-\psi))\subset L^{1}(0,T;H)$. Note that $\|(\bar{B_{0}}w_{n}+g_{n})\psi_{n}\|_{H}\leq|\!|\!|\bar{B_{0}}w_{n}+g_{n}|\!|\!|_{H}\|\psi_{n}\|_{U}$ with $(\psi_{n})$ bounded in $L^{2}(0,T;U)$ and similar for ${b}(\psi_{n}-\psi)$. Thus by the Cauchy–Schwarz inequality, if we show that $(\bar{B_{0}}w_{n})$ and $(g_{n})$ are bounded in $L^{2}(0,T;\mathcal{L}_{2}(U,H))$, then boundedness of $(\beta_{n})\subset L^{1}(0,T;H)$ follows (${b}\in L^{2}(0,T;\mathcal{L}_{2}(U,H))$ does not depend on $n$). By symmetry in Assumption 2.2(3), $\bar{B_{0}}w_{n}$ and $g_{n}$ can be estimated in the same way as $\bar{A_{0}}w_{n}\coloneqq A_{0}(u^{\psi})w_{n}$ and $f_{n}\coloneqq(A_{0}(u^{\psi})-A_{0}(u^{\psi_{n}}))u^{\psi_{n}}+F(u^{\psi_{n}})-F(u^{\psi})$, respectively. We provide the estimates for the latter here. By Lemma 3.8(ii)(iii), $\|A_{0}(u^{\psi})w_{n}\|_{L^{2}(0,T;V^{*})}\leq C_{N,T}(1+N)N<\infty$ and $\|(A_{0}(u^{\psi})-A_{0}(u^{\psi_{n}}))u^{\psi_{n}}\|_{L^{2}(0,T;V^{*})}\leq C_{N,T}N^{2}<\infty$. Furthermore, Lemma 3.8(iv) gives $\|F(u^{\psi_{n}})\|_{L^{2}(0,T;V^{*})}\leq\tilde{C}_{N,T}(1+N)<\infty$. Finally, $F(u^{\psi})\in L^{2}(0,T;V^{*})$ does not depend on $n$. We conclude that $(-\bar{A_{0}}w_{n}+f_{n})$ is bounded in $L^{2}(0,T;V^{*})$ and by the considerations above, $(\beta_{n})$ is bounded in $L^{1}(0,T;H)$. Lemma 4.8 thus yields $\lim_{n\to\infty}\sup_{t\in[0,T]}|I_{3}^{n}(t)|=0$ and (4.24) gives $w_{n}\to 0$ in $\mathrm{MR}(0,T)$. ∎
@@
Proposition 4.9 also ensures measurability of the map $\mathcal{G}^{0}\colon C([0,T];U_{1})\to\mathrm{MR}(0,T)$ defined by (4.4), as required in Theorem 4.5. Note that $\{\int_{0}^{\cdot}\psi(s)\,\mathrm{d}s:\psi\in L^{2}(0,T;U)\}=\{v\in W^{1,2}(0,T;U):v(0)=0\}\eqqcolon W^{1,2}_{0}$. By Sobolev embedding [HNVWvolume3, Corollary L.4.6], $W^{1,2}_{0}$ embeds continuously into $C([0,T];U)$ ($W^{1,2}_{0}$ is a closed subspace of $W^{1,2}(0,T;U)$). Hence Kuratowski’s theorem [kechris95, Th. 15.1] gives $\mathcal{B}(W^{1,2}_{0})\subset\mathcal{B}(C([0,T];U))$. Moreover, $\gamma\colon W^{1,2}_{0}\to\mathrm{MR}(0,T)\colon\int_{0}^{\cdot}\psi(s)\,\mathrm{d}s\mapsto u^{\psi}$ is continuous, since $\int_{0}^{\cdot}\psi_{n}(s)\,\mathrm{d}s\to\int_{0}^{\cdot}\psi(s)\,\mathrm{d}s$ in $W^{1,2}_{0}$ implies $\psi_{n}\to\psi$ in $L^{2}(0,T;U)$, and $L^{2}(0,T;U)\to\mathrm{MR}(0,T)\colon\psi\mapsto u^{\psi}$ is norm-continuous since it is weakly sequentially continuous by Proposition 4.9. It follows that for $E\in\mathcal{B}(\mathrm{MR}(0,T))$, we have
@@ 4.3. Stochastic continuity criterion
It remains to verify the stochastic continuity criterion (iii) of Theorem 4.5. Before we prove that (iii) is satisfied, we first derive some stochastic bounds which we will later apply to $X^{\varepsilon}\coloneqq\mathcal{G}^{\varepsilon}\big{(}\tilde{W}_{1}(\cdot)+\frac{1}{\sqrt{\varepsilon}}\int_{0}^{\cdot}\Psi^{\varepsilon}(s)\,\mathrm{d}s\big{)}$. In the next lemma we use a stochastic Gronwall lemma as in [AV22variational] to avoid further growth bound assumptions on $B$.
@@
$\begin{cases}\sup_{\varepsilon\in(0,\frac{1}{2})}\mathbb{P}(\|X^{\varepsilon}\|_{\mathrm{MR}(0,T)}>\gamma)\leq\frac{C}{\gamma^{2}},\\ \sup_{\varepsilon\in(0,\frac{1}{2})}\mathbb{P}(\|B(\cdot,X^{\varepsilon}(\cdot))\|_{L^{2}(0,T;\mathcal{L}_{2}(U,H))}>\gamma)<\frac{C}{\gamma^{2}}.\end{cases}$
@@
$\qquad\qquad+2\sqrt{\varepsilon}\int_{0}^{t}\langle X^{\varepsilon}(s),B(s,X^{\varepsilon}(s))\,\mathrm{d}W(s)\rangle+\varepsilon\int_{0}^{t}|\!|\!|B(s,X^{\varepsilon}(s))|\!|\!|_{H}^{2}\,\mathrm{d}s$
$\leq 2\int_{0}^{t}-\frac{1}{2}|\!|\!|B(s,X^{\varepsilon}(s))|\!|\!|_{H}^{2}-\theta\|X^{\varepsilon}(s)\|_{V}^{2}+M\|X^{\varepsilon}(s)\|_{H}^{2}+|\phi(s)|^{2}\,\mathrm{d}s$
@@
$\leq-\int_{0}^{t}|\!|\!|B(s,X^{\varepsilon}(s))|\!|\!|_{H}^{2}\,\mathrm{d}s-2\theta\|X^{\varepsilon}\|_{L^{2}(0,t;V)}^{2}+\int_{0}^{t}2M\|X^{\varepsilon}(s)\|_{H}^{2}\,\mathrm{d}s+2\|\phi\|_{L^{2}(0,t)}^{2}$
$\qquad+2\int_{0}^{t}\frac{1}{8}|\!|\!|B(s,X^{\varepsilon}(s))|\!|\!|_{H}^{2}+2\|\Psi^{\varepsilon}(s)\|_{U}^{2}\|X^{\varepsilon}(s)\|_{H}^{2}\,\mathrm{d}s$
@@
Now the stochastic Gronwall inequality [geiss24, Cor. 5.4b), (50)] (with $R\coloneqq 2MT+4K^{2}$) gives
@@
$\mathbb{P}(\|X^{\varepsilon}\|_{\mathrm{MR}(0,T)}^{2}>\gamma)<\frac{C}{\gamma},\quad\mathbb{P}(\|B(\cdot,X^{\varepsilon}(\cdot))\|_{L^{2}(0,T;\mathcal{L}_{2}(U,H))}^{2}>\gamma)<\frac{C}{\gamma}.$
Consequently, we have $\mathbb{P}(\|X^{\varepsilon}\|_{\mathrm{MR}(0,T)}>\gamma)=\mathbb{P}(\|X^{\varepsilon}\|_{\mathrm{MR}(0,T)}^{2}>\gamma^{2})\leq\frac{C}{\gamma^{2}}$ and in the same way, $\mathbb{P}(\|B(\cdot,X^{\varepsilon}(\cdot))\|_{L^{2}(0,T;\mathcal{L}_{2}(U,H))}>\gamma)<\frac{C}{\gamma^{2}}$, uniformly in $\varepsilon\in(0,\frac{1}{2})$. ∎
@@
$\mathbb{P}(\|X^{\varepsilon}-u^{\varepsilon}\|_{\mathrm{MR}(0,T)}>\gamma)$ ($\leq\mathbb{P}(\{\|X^{\varepsilon}-u^{\varepsilon}\|_{\mathrm{MR}(0,T)}>\gamma\}\cap E_{n,\varepsilon})+\mathbb{P}(E_{n,\varepsilon}^{c})$
($\leq\mathbb{P}(\{\|X^{\varepsilon}-u^{\varepsilon}\|_{\mathrm{MR}(0,T)}>\gamma\}\cap E_{n,\varepsilon})+\frac{C}{n^{2}}.$
@@
($\mathbb{P}(\{\|X^{\varepsilon}-u^{\varepsilon}\|_{\mathrm{MR}(0,T)}^{2}>\delta\}\cap E_{n,\varepsilon})\leq\sum_{i=3}^{4}\mathbb{P}(\{\sup_{t\in[0,T]}|{I}_{i}^{\varepsilon}(t)|>\frac{\delta}{2c_{n}}\}\cap E_{n,\varepsilon}).$
@@
$\qquad+\int_{0}^{t}\frac{1}{2}|\!|\!|B_{0}(s,u^{\varepsilon}(s))(X^{\varepsilon}(s)-u^{\varepsilon}(s))|\!|\!|_{H}^{2}+\frac{1}{2}\|\Psi^{\varepsilon}(s)\|_{U}^{2}\|X^{\varepsilon}(s)-u^{\varepsilon}(s)\|_{H}^{2}\,\mathrm{d}s$
@@
$\qquad+\int_{0}^{t}|\!|\!|\big{(}B_{0}(s,X^{\varepsilon}(s))-B_{0}(s,u^{\varepsilon}(s))\big{)}X^{\varepsilon}(s)|\!|\!|_{H}\|\Psi^{\varepsilon}(s)\|_{U}\|X^{\varepsilon}(s)-u^{\varepsilon}(s)\|_{H}\,\mathrm{d}s$
$\qquad+\int_{0}^{t}|\!|\!|G(X^{\varepsilon}(s))-G(u^{\varepsilon}(s))|\!|\!|_{H}\|\Psi^{\varepsilon}(s)\|_{U}\|X^{\varepsilon}(s)-u^{\varepsilon}(s)\|_{H}\,\mathrm{d}s$
@@
$\qquad+\int_{0}^{t}\frac{1}{2}|\!|\!|\big{(}B_{0}(s,X^{\varepsilon}(s))-B_{0}(s,u^{\varepsilon}(s))\big{)}X^{\varepsilon}(s)|\!|\!|_{H}^{2}+\frac{1}{2}\|\Psi^{\varepsilon}(s)\|_{U}^{2}\|X^{\varepsilon}(s)-u^{\varepsilon}(s)\|_{H}^{2}\,\mathrm{d}s$
$\qquad+\frac{1}{2}\int_{0}^{t}|\!|\!|G(X^{\varepsilon}(s))-G(u^{\varepsilon}(s))|\!|\!|_{H}^{2}+\frac{1}{2}\|\Psi^{\varepsilon}(s)\|_{U}^{2}\|X^{\varepsilon}(s)-u^{\varepsilon}(s)\|_{H}^{2}\,\mathrm{d}s$
@@
It remains to prove (4.31). Note that ${I}_{4}^{\varepsilon}$ is a continuous local martingale (using Lemma 4.11) starting at zero, with $[{I}_{4}^{\varepsilon}](T)=\int_{0}^{T}\varepsilon\|\langle X^{\varepsilon}(s)-u^{\varepsilon}(s),B(s,X^{\varepsilon}(s))(\cdot)\rangle\|_{\mathcal{L}_{2}(U,{\mathbb{R}})}^{2}\,\mathrm{d}s$, where $[I_{4}^{\varepsilon}]$ denotes the quadratic variation. Thus, by [kallenberg21, Prop. 18.6], (4.31) is equivalent to
@@
$\leq\mathbb{P}\Big{(}\int_{0}^{T}\|X^{\varepsilon}(s)-u^{\varepsilon}(s)\|_{H}^{2}|\!|\!|B(s,X^{\varepsilon}(s))|\!|\!|_{H}^{2}\,\mathrm{d}s>{\delta}{\varepsilon}^{-1}\Big{)}$
$\leq\mathbb{P}\Big{(}\|X^{\varepsilon}-u^{\varepsilon}\|_{C([0,T];H)}^{2}\int_{0}^{T}|\!|\!|B(s,X^{\varepsilon}(s))|\!|\!|_{H}^{2}\,\mathrm{d}s>{\delta}{\varepsilon}^{-1}\Big{)}$

### v5

@@ 4.1. Weak convergence approach
if and only if $Y^{\varepsilon}$ is a strong solution in the sense of [undefe, Def. 3.2] to (4.1) with $x$ replaced by $\xi$. This is a mere consequence of (A.7) and the fact that $B(\cdot,Y(\cdot))\in L^{2}(([0,T]\times\Omega,\mathcal{P},\lambda\times\mathbb{P});\mathcal{L}_{2}(U,H))\subset\mathcal{N}(0,T)$ for any $Y\in\mathrm{MR}(0,T)$. By [undefe, Th. 3.5], (4.1) has a unique strong solution $Y^{\varepsilon}$, also when $x$ is replaced by random initial data $\xi$. Thus we have pathwise uniqueness in the sense of [undefam, Def. 1.7] and we have existence of a.s. $\mathrm{MR}(0,T)$-valued weak solutions. Now fix $x\in H$ and $\varepsilon>0$ and let $Y^{\varepsilon}$ be the unique strong solution to (4.1). By [undefam, Th. 2.1, Def. 1.9(2), Def. 1.8] there exists a measurable map $\mathcal{G}^{\varepsilon}\colon C([0,T];U_{1})\to\mathrm{MR}(0,T)$ such that a.s. $Y^{\varepsilon}=\mathcal{G}^{\varepsilon}(\tilde{W}_{1}(\cdot))$.
@@
We have $\mathbb{E}[\exp(\frac{1}{2}\|-\frac{1}{\sqrt{\varepsilon}}\Psi^{\varepsilon}\|_{L^{2}(0,T;U)}^{2})]\leq\exp(\frac{K^{2}}{2\varepsilon})<\infty$, so by Novikov’s condition [undefw, Prop. 5.12],
@@
[…]
$\hat{\mathbb{P}}\coloneqq\exp\left(-\frac{1}{\sqrt{\varepsilon}}\int_{0}^{T}\langle\Psi^{\varepsilon}(s),\,\mathrm{d}\tilde{W}(s)\rangle_{U}-\frac{1}{2\varepsilon}\|\Psi^{\varepsilon}\|_{L^{2}(0,T;U)}^{2}\right)\,\mathrm{d}\mathbb{P}.$
@@
[…]
@@
where the last equality follows from the definition of $\hat{\beta}_{k}$ and (A.6). Let $\hat{\mathcal{N}}(0,T)$ denote the stochastically integrable processes with respect to $\hat{\mathcal{W}}$ and $\hat{W}$ on $(\Omega,\mathcal{F},\hat{\mathbb{P}},(\mathcal{F}_{t})_{t\geq 0})$, i.e. (A.3) with $\mathbb{P}$ replaced by $\hat{\mathbb{P}}$. Note that $\hat{\mathcal{N}}(0,T)=\mathcal{N}(0,T)$, since $\mathbb{P}\ll\hat{\mathbb{P}}\ll\mathbb{P}$. Thus, Proposition A.4 gives $\int_{0}^{t}\Phi(s)\,\mathrm{d}\hat{\mathcal{W}}(s)=\int_{0}^{t}\Phi(s)\,\mathrm{d}\hat{W}(s)$ $^\hat{\mathbb{P}}$-a.s. for all $\Phi\in\mathcal{N}(0,T)$ and $t\in[0,T]$. Therefore, combined with (4.6), ${X}^{\varepsilon}$ satisfies $^\hat{\mathbb{P}}$-a.s. (hence $\mathbb{P}$-a.s.) in $V^{*}$:
@@
In the last line we used that $\int_{0}^{t}\Phi(s)\,\mathrm{d}\hat{\mathcal{W}}(s)=\int_{0}^{t}\Phi(s)\,\mathrm{d}{W}(s)+\frac{1}{\sqrt{\varepsilon}}\int_{0}^{t}\Phi(s)\Psi^{\varepsilon}(s)\,\mathrm{d}s$ for $\Phi\in\mathcal{N}(0,T)$ and $t\in[0,T]$. For $\Phi=\mathbbm{1}_{A\times(t_{1},t_{2}]}\otimes(u\otimes x)$ with $0\leq t_{1}<t_{2}\leq T$, $A\in\mathcal{F}_{t_{1}}$, $u\in U$, $x\in H$, the identity follows from (4.7) and the definition of the stochastic integral for elementary processes [undefad, p. 305]. By linearity and continuity of the integrals and by a density argument and localization, the identity extends for $\Phi\in\mathcal{N}(0,T)$. This finishes the proof of the last claim of the lemma. ∎
@@ 4.2. Weakly continuous dependence in the skeleton equation
In this subsection we prove that condition (ii) of Theorem 4.5 is satisfied. This will be achieved in the upcoming Proposition 4.9. Its proof was inspired by [undefs, Th. 3.2]. Using an additional approximation by Bochner-simple functions, we can omit the time(-Hölder) regularity assumptions on $B$ of [undefs, (H5)].
@@
To estimate further, we use an argument inspired by [undefs, Lem. 3.3]. Note that $w_{n}(0)=0$ and
@@
For any $t\in[\delta,T]$, we can apply the chain rule (A.2) to $v_{n}^{t,\delta}(\cdot)\coloneqq w_{n}(\cdot)-w_{n}(t_{\delta})=\int_{t_{\delta}}^{\cdot}\alpha_{n}(s)\,\mathrm{d}s$ on $t_{\delta},T]$ and obtain for all $\tilde{t}\in[t_{\delta},T]$:
@@
[…]
@@
[…]
@@
[…]
@@
Proposition 4.9 also ensures measurability of the map $\mathcal{G}^{0}\colon C([0,T];U_{1})\to\mathrm{MR}(0,T)$ defined by (4.4), as required in Theorem 4.5. Note that $\{\int_{0}^{\cdot}\psi(s)\,\mathrm{d}s:\psi\in L^{2}(0,T;U)\}=\{v\in W^{1,2}(0,T;U):v(0)=0\}\eqqcolon W^{1,2}_{0}$. By Sobolev embedding [undefu, Corollary L.4.6], $W^{1,2}_{0}$ embeds continuously into $C([0,T];U)$ ($W^{1,2}_{0}$ is a closed subspace of $W^{1,2}(0,T;U)$). Hence Kuratowski’s theorem [undefx, Th. 15.1] gives $\mathcal{B}(W^{1,2}_{0})\subset\mathcal{B}(C([0,T];U))$. Moreover, $\gamma\colon W^{1,2}_{0}\to\mathrm{MR}(0,T)\colon\int_{0}^{\cdot}\psi(s)\,\mathrm{d}s\mapsto u^{\psi}$ is continuous, since $\int_{0}^{\cdot}\psi_{n}(s)\,\mathrm{d}s\to\int_{0}^{\cdot}\psi(s)\,\mathrm{d}s$ in $W^{1,2}_{0}$ implies $\psi_{n}\to\psi$ in $L^{2}(0,T;U)$, and $L^{2}(0,T;U)\to\mathrm{MR}(0,T)\colon\psi\mapsto u^{\psi}$ is norm-continuous since it is weakly sequentially continuous by Proposition 4.9. It follows that for $E\in\mathcal{B}(\mathrm{MR}(0,T))$, we have
@@ 4.3. Stochastic continuity criterion
It remains to verify the stochastic continuity criterion (iii) of Theorem 4.5. Before we prove that (iii) is satisfied, we first derive some stochastic bounds which we will later apply to $X^{\varepsilon}\coloneqq\mathcal{G}^{\varepsilon}\big(\tilde{W}_{1}(\cdot)+\frac{1}{\sqrt{\varepsilon}}\int_{0}^{\cdot}\Psi^{\varepsilon}(s)\,\mathrm{d}s\big)$. In the next lemma we use a stochastic Gronwall lemma as in [undefe] to avoid further growth bound assumptions on $B$.
@@
$\begin{cases}\sup_{\varepsilon\in(0,\frac{1}{2})}\mathbb{P}(\|X^{\varepsilon}\|_{\mathrm{MR}(0,T)}>\gamma)\leq\frac{C}{\gamma^{2}},\\ \sup_{\varepsilon\in(0,\frac{1}{2})}\mathbb{P}(\|B(\cdot,X^{\varepsilon}(\cdot))\|_{L^{2}(0,T;\mathcal{L}_{2}(U,H))}>\gamma)\leq\frac{C}{\gamma^{2}}.\end{cases}$
@@
[…]
@@
[…]
@@
Now the stochastic Gronwall inequality [undefr, Cor. 5.4b), (50)] (with $R\coloneqq 2MT+4K^{2}$) gives
@@
$\mathbb{P}(\|X^{\varepsilon}\|_{\mathrm{MR}(0,T)}^{2}>\gamma)\leq\frac{C}{\gamma},\quad\mathbb{P}(\|B(\cdot,X^{\varepsilon}(\cdot))\|_{L^{2}(0,T;\mathcal{L}_{2}(U,H))}^{2}>\gamma)\leq\frac{C}{\gamma}.$
Consequently, we have $\mathbb{P}(\|X^{\varepsilon}\|_{\mathrm{MR}(0,T)}>\gamma)=\mathbb{P}(\|X^{\varepsilon}\|_{\mathrm{MR}(0,T)}^{2}>\gamma^{2})\leq\frac{C}{\gamma^{2}}$ and in the same way, $\mathbb{P}(\|B(\cdot,X^{\varepsilon}(\cdot))\|_{L^{2}(0,T;\mathcal{L}_{2}(U,H))}>\gamma)\leq\frac{C}{\gamma^{2}}$, uniformly in $\varepsilon\in(0,\frac{1}{2})$. ∎
@@
$\mathbb{P}(\|X^{\varepsilon}-u^{\varepsilon}\|_{\mathrm{MR}(0,T)}>\gamma)$ $\leq\mathbb{P}(\{\|X^{\varepsilon}-u^{\varepsilon}\|_{\mathrm{MR}(0,T)}>\gamma\}\cap E_{n,\varepsilon})+\mathbb{P}(E_{n,\varepsilon}^{c})$
$\leq\mathbb{P}(\{\|X^{\varepsilon}-u^{\varepsilon}\|_{\mathrm{MR}(0,T)}>\gamma\}\cap E_{n,\varepsilon})+\frac{C}{n^{2}}.$
@@
$\mathbb{P}(\{\|X^{\varepsilon}-u^{\varepsilon}\|_{\mathrm{MR}(0,T)}^{2}>\delta\}\cap E_{n,\varepsilon})\leq\sum_{i=3}^{4}\mathbb{P}(\{\sup_{t\in[0,T]}|{I}_{i}^{\varepsilon}(t)|>\frac{\delta}{2c_{n}}\}\cap E_{n,\varepsilon}).$
@@
[…]
@@
[…]
@@
[…]
@@
It remains to prove (4.31). Note that ${I}_{4}^{\varepsilon}$ is a continuous local martingale (using Lemma 4.11) starting at zero, with ${I}_{4}^{\varepsilon}](T)=\int_{0}^{T}\varepsilon\|\langle X^{\varepsilon}(s)-u^{\varepsilon}(s),B(s,X^{\varepsilon}(s))(\cdot)\rangle\|_{\mathcal{L}_{2}(U,{\mathbb{R}})}^{2}\,\mathrm{d}s$, where $I_{4}^{\varepsilon}]$ denotes the quadratic variation. Thus, by [undefv, Prop. 18.6], (4.31) is equivalent to
@@
[…]